Percent composition of oxygen in kclo3
1
Expert's answer
2020-01-17T07:01:45-0500

First, we calculate the molar mass of oxygen in this compound.


M (O) = 16 * 3 = 48 g/mol.


Next, we calculate the molar mass of the entire compound. Write down the solution.


M (KClO3) = 39 + 35,5 + 16 * 3 = 122,5 g / mol.


Next, we find the mass fraction of oxygen in this compound.


W = 48/122,5 * 100 % = 39,1 %.


This means that the mass fraction of oxygen is 39.1 %